Use Multivariate Transformations 7. (10 marks) The joint density of X and Y is given as f(x,y)=10, else a) (5 marks) Find the joint probability density function of = X / y and V = X. b) (5 marks) Using your result in (a), find the marginal density of V.
Answer #1

We will use Jacobian method to find the joint probability density function of U and V. It is given as

f(u, v) = f(x,y) |J|

where,

x, y are the solutions of the equation u = x/y and v = xy

and

J = egin{vmatrix} rac{partial x}{partial u} & rac{partial x}{partial v} rac{partial y}{partial u} & rac{partial y}{partial v} end{vmatrix}

For 0 < y < 1 and 0 < x < 1

u can range from 0 to infty and v can range from 0 to 1

20 t2 2 else du and lz

20U